A stretched string fixed at each end has a mass of 48.0 g and a length of 7.00 m. The tension in the string is 55.0 N
a) Determine the positions of the nodesand antinodes for the third harmonic (Enter your answers from smallest to largest distance from one end of the string.)
b) What is the vibration frequency for this harmonic?
Answer #1

Given the length of the wire = 7 m

For the third harmonic frequency, there are 3 loops in the wave

Thus the length of each loop, λ/2 = 7/3

So the wave length, λ = 4.67 m

a) As there are three loops in the third harmonic,

The nodes are occur at the distances 0, λ/2, λ, 3λ/2

i.e. the nodes are occur at 0, 2.33m, 4.67m, 7m.

The antinodes will occur halfway between each pair of adjacent nodes so that their positions will be at, 1.165m, 3.495m and 5.825m successively from one end.

b) The vibrational frequency is given by the formula

       f = v/λ

linear density, μ = m/l = 0.048/7 = 0.00686 kg/m

and Tension force, T = 55 N

velocity of the wave, v = (T/μ)^(1/2) = 89.56 m/s

f = 89.56/4.67 = 19.18 Hz

Kindly upvote :)